# JiraiBench: A Bilingual Benchmark for Evaluating Large Language Models’ Detection of Human Self-Destructive Behavior Content in Jirai Community

### v2

@@ Abstract
This paper introduces JiraiBench, the first bilingual benchmark for evaluating large language models’ effectiveness in detecting self-destructive content across Chinese and Japanese social media communities. Focusing on the transnational ”Jirai” (landmine) online subculture that encompasses multiple forms of self-destructive behaviors including drug overdose, eating disorders, and self-harm, we present a comprehensive evaluation framework incorporating both linguistic and cultural dimensions. Our dataset—comprising 10,419 Chinese posts and 5,000 Japanese posts—features multidimensional annotation along three behavioral categories with substantial inter-annotator agreement. Experimental evaluations across four state-of-the-art models reveal significant performance variations based on instructional language, with Japanese prompts unexpectedly outperforming Chinese prompts when processing Chinese content. This emergent cross-cultural transfer suggests that cultural proximity can sometimes outweigh linguistic similarity in detection tasks. Cross-lingual transfer experiments with fine-tuned models further demonstrate the potential for knowledge transfer between these language systems without explicit target language training. These findings highlight the need for culturally informed approaches to multilingual content moderation and provide empirical evidence for the importance of cultural context in developing more effective detection systems for vulnerable online communities.
Disclaimer: This paper describes human self-destructive content and potentially harmful behaviors including drug overdoese, eating disorders, and self-harming actions that may be disturbing to some readers.
Jirai Kei ”Jirai Kei” (in English: landmine style; in Chinese: 地雷系; in Japanese: じらいけい) refers to a transnational social media subculture that emerged in Japan during the early 2020s, gaining particular prominence during pandemic restrictions. The term literally translates to ”landmine style”, metaphorically representing psychological vulnerabilities that might ”explode” when triggered. This community is characterized by individuals who publicly express various forms of psychological distress and self-destructive tendencies across social media platforms. Jirai communities exhibit distinctive communication patterns, including specialized terminology and coded language that conveys emotional vulnerability, particularly related to eating disorders, self-harm behaviors, and substance misuse. This phenomenon spans both Japanese and Chinese online spaces, creating interconnected networks where participants share experiences of psychological distress through culturally specific expressions.
@@ 1 Introduction
Human self-destructive behaviors constitute a profound public health challenge (Baumeister & Scher, 1988). These behaviors manifest as deliberate acts of self-harm that can range from subtle to severe, including eating disorders (ED) characterized by extreme dietary restriction and purging, self-harm (SH) by cutting or burning, and drug overdose (OD) at levels that risk permanent harm or death. What makes these behaviors particularly concerning is their tendency to operate beneath the surface of conventional detection systems, often interwoven with complex mental health conditions such as depression, anxiety, post-traumatic stress disorder, or personality disorders (Van der Kolk et al., 1991; Firestone & Seiden, 1990). The digital era has further complicated this landscape, as online communities have emerged that sometimes normalize, reinforce, or provide methods for self-destructive behaviors, creating new challenges for intervention and prevention efforts.
Previous research efforts in detecting self-destructive behaviors online have made notable contributions but remain limited in scope and applicability. Most studies have focused narrowly on individual categories of concerning behaviors, examining eating disorders (ED) (Wang et al., 2017a; Moessner et al., 2018; Chancellor et al., 2016a), self-harm (SH) (Wang et al., 2017b; Un Nisa & Muhammad, 2021; Ragheb et al., 2021), or drug overdose (OD) (Fisher et al., 2023; Nasralah et al., 2020; Phan et al., 2017; Fan et al., 2017) as isolated phenomena. These approaches have yielded domain-specific insights and specialized detection techniques, establishing methodological foundations for computational analysis of psychological distress signals in digital contexts. Additionally, community-based detection approaches (Tébar & Gopalan, 2021; Wang et al., 2017a; Chancellor et al., 2016b) have demonstrated promising results by analyzing linguistic patterns within specific online groups, though these implementations remain relatively sparse in the literature.
However, these existing approaches exhibit significant limitations that constrain their effectiveness in real-world applications. First, the predominant focus on single behavioral categories fails to address the complex comorbidities and overlapping patterns frequently observed in at-risk individuals, who often exhibit multiple forms of self-destructive behaviors simultaneously. Second, the overwhelming concentration on English-language content (Scherr et al., 2020; Tébar & Gopalan, 2021; Sixto-García & Duarte-Melo, 2020) severely restricts applicability in multilingual environments where vulnerable communities communicate using culturally-specific expressions and coded terminology, creating substantial blind spots in global detection systems. Furthermore, the limited adoption of community-contextualized approaches neglects crucial social dynamics through which self-destructive ideation becomes normalized and reinforced, significantly reducing model sensitivity when confronted with the nuanced, evolving discourse of real-world at-risk communities. These constraints collectively highlight the need for more comprehensive, multilingual, and contextually-aware detection frameworks that can effectively identify concerning content across diverse cultural and linguistic contexts.
To address existing gaps in research and intervention, our study examines the Jirai (in English: landmine; in Chinese:地雷; in Japanese: じらい) community [^1] https://aesthetics.fandom.com/wiki/Jirai_Kei—a transnational social media phenomenon spanning Chinese and Japanese online spaces that encompasses multiple forms of self-destructive behaviors simultaneously. This community represents an ideal case study for developing JiraiBench, as it presents unique challenges for content moderation systems and mental health interventions. What makes the Jirai community particularly valuable for our research is its simultaneous manifestation of self-destructive behaviors such as substance abuse, eating disorders, and self-harm—precisely the categories targeted by our annotation framework. This stands in contrast to previous research that typically focuses on single behavioral domains. Moreover, the Jirai community poses complex challenges due to its cross-border nature across differing regulatory environments, the use of coded language to evade standard monitoring algorithms, internal reinforcement mechanisms that intensify harmful behaviors, and the persistent stigma surrounding mental health issues in both Chinese and Japanese societies. These factors often divert vulnerable individuals toward alternative peer networks rather than seeking professional help (Niramaya & Filia, 2023).
By analyzing this bilingual community through a contextually aware approach, we demonstrate how language-specific nuances and community dynamics must be incorporated into detection systems—addressing a significant limitation of English-centric, behavior-isolated models that lack the cultural and linguistic sensitivity required for effective intervention in diverse global contexts.
@@
1) We conducted comprehensive experiments in 4 state-of-the-art LLMs using 2 baseline configurations, systematically evaluating performance in both target languages with prompts in Chinese, Japanese, and English, revealing unexpected patterns of cross-cultural transfer.
2) We uncover a noteworthy emergent phenomenon in which Japanese instruction prompts consistently outperform Chinese prompts when processing Chinese content, suggesting important linguistic-cultural bridges in transnational content moderation tasks.
3) We investigate cross-lingual transfer capabilities by fine-tuning Qwen2.5 7B on Chinese data and demonstrate significant performance improvements on both source and target languages, particularly by improving detection of self-destructive behaviors in Japanese content without explicit Japanese training data.
@@ 2.1 Social Media and Mental Health
Natural language processing (NLP) has long been a pivotal tool for researchers who studied mental health problems through textual data, ranging from conventional clinical records to contemporary digital platforms content. Social media was regarded as a valuable resource to understand users mental dynamics, with studies showing that such content carries detectable signals of mental disorders and that online activities can reflect individuals’ pyschological states (Paul & Dredze, 2011; Coppersmith et al., 2014).
@@
Among the considerable amount of NLP-based mental health studies, depression is the most discussed one due to its high prevalence and broad implications. It has been well established that individuals’ mental states can be conveyed through language use (Chancellor & De Choudhury, 2020), one representative example of which is that depressed people tend to use first-person pronouns and negative words with a higher frequency (De Choudhury et al., 2013). The boundary of NLP research in mental health was also extended to other types as mental illness has become a shared, concerning issue for contemporary human beings. Post-Traumatic Stress Disorder (PTSD) (Murarka et al., 2021), Attention Deficit Hyperactivity Disorder (ADHD) (Guntuku et al., 2023), and other general problems have all been examined via language cues, such as references to trauma, heightened expressions of fear, expressions of worry, and rumination. These linguistic markers are usually emblematic of negative psychological conditions, exposing critical insights into how mental states manifest in languages (Teodorescu et al., 2023).
@@
The evolution of NLP nurtured the growing research interest in this domain. As this field matured in mid-2010s, research focus shifted from domain-specific lexical analysis and correlation statistics to machine learning-based predictive models. Behavioral features, such as linguistic style, posting frequency, and social network attributes were engineered and then fed into classifiers, such as logistic regression (LR) and support vector machine (SVM), to detect relevant user-generated content (De Choudhury et al., 2013; Tsugawa et al., 2015).
@@
With the introduction of transformer, transformer-based models revolutionized the research landscape in this area, bringing remarkable improvements in the detection task and outperformed earlier deep learning models (Jiang et al., 2020; Matero et al., 2019). In particular, specialized models like MentalBERT that were trained in a mental health-specific corpus have demonstrated an improved capability of identifying edged language patterns in people’s mental health disclosures (Ji et al., 2022).
@@
Although the prior methods have shown promising results, several key challenges persist. One recurring issue is data scarcity and quality. Clinical diagnostic data is normally private and scarce, while datasets populated with online posts were in the doubt of construct validity—the lack of a standardized assessment of each post not only introduces mislabeling risk but also complicates the process of results evaluation. Another pivotal issue lies in the peculiar, intricate linguistic patterns in posts involved with negative mental conditions. Usually, distress was expressed in a figurative or implicit way, where creative metaphors, sarcasm, and irony were communicated as an encoded language to describe the feelings (Coppersmith et al., 2014; Mendes & Caseli, 2024). Additionally, most research were established on English dataset with far fewer in other languages such as Chinese and Japanese (Cao et al., 2025) and models therein were trained on corpus constructed from single platform or a homogeneous demographic, consequently limiting models applicability to more diverse populations.
2.2 NLP for Self-Destructive Behaviors
@@
Besides the shared attributes, targeting self-destructive behaviors differs remarkably from the task on general mental illness. Self-destructive behaviors that might incur detrimental and even life-threatening outcomes necessitate more accurate detection for taking preventive actions, while their unique linguistic patterns confront researchers with more challenges.
@@
Flagging users on social media who reveal a tendency toward non-suicidal SH acts and suicidal ideation is a vital yet sensitive application of NLP. Studies reveal that the combination of linguistic signals from content, such as hopelessness, self-hate, and past suicidal attempts, coupled with online behavioral patterns greatly improve the models performance—and it usually gets better when instilled with domain knowledge (Zhang et al., 2024; Ng et al., 2023; Choudhury & Kıcıman, 2017). Some works focus specifically on SH that might include mentions of self-cutting or injury (Cliffe et al., 2021), where the incorporation of psychological lexicons helps models identify depressive expressions as warning sign.
@@
Likewise, ED detection is another intractable task because the involved individuals barely explicitly state their conditions. Instead, such tendencies manifest through obsessive discussions of weight, food, calories, and body images, which makes the ingestion of specific knowledge essential to ensure models performance on the classification task accordingly (Chan et al., 2022).
@@
Social media-based OD study is also one of the concerning areas that NLP is widely applied to (Wright et al., 2021). One unique challenge specific to this domain is the language ambiguity of drug references and the fast-evolving drug landscape, further complicating the construction of datasets—from identifying emerging drugs through obscure drug references to discerning the intent behind those drug-included posts. Annotating such data also requires expert assistance in analyzing if the post is truly indicative of substance abuse rather than casual innocuous mentioning.
Most recently, the surge of LLMs, such as GPT-4 (et el, 2024), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023), have create new avenues for tasks in mental healthcare. LLMs demonstrate strong capabilities in distinguishing user-generated content related to SD behaviors and catching on the worsening psychological conditions of sufferers, such as disordered eating pathology,substance abuse, and suicidal ideation—even in zero-shot or few-shot settings (Stade et al., 2024). Based on their supremacy in understanding the linguistic nuances of the intense content of information on social media, such models can identify subtextual signals indicative of mental distress, enabling immediate interventions.
However, LLMs-based approach for self-destructive content detection is faced with generalizability obstacles that span across platforms, cultures, and languages thereof. Most mainstream LLMs were primarily trained on English-centric corpus, whereas materials in other languages only take a fairly small proportion of that. Even these state-of-the-art models struggle while adapting to multilingual environments featuring different languages and cultures—this task becomes trickier given the community-based, culture-dependent expressions. While LLMs excel at contextual understanding, the linguistic peculiarities varying by users across platforms can imperil the models performance, especially when expressions are intentionally obfuscated to evade content moderation.
@@ 3 Dataset Construction
In this section, we describe the construction of JiraiBench dataset. We start with introducing the data collection and filtering process, then turning to the annotation.
@@ 3.1 Data Collection and Cleaning
For the Chinese dataset, we implemented a keyword-based search approach to gather relevant posts from Sina Weibo [^2] https://weibo.com/, China’s equivalent of X (formerly Twitter). Using specialized lexicons developed by domain experts in both languages for OD, ED, and SH, we identified posts that discuss self-destructive behaviors within the community. This process yielded a comprehensive Chinese dataset containing 10,419 potentially concerning posts. The lexicons were meticulously curated to include both explicit and implicit terminology commonly associated with each behavior type. Similarly, we applied the same methodological framework to construct the Japanese dataset, ensuring consistency between the collections. From X [^3] https://x.com/, we collected 5,000 tweets, with varying associations to the three behaviors of interest. Both datasets underwent an identical data cleansing pipeline to remove irrelevant content, including advertisements, duplicate entries, excessively short posts lacking meaningful information, and other noise. In accordance with platform policies and to maintain user privacy, we eliminated all usernames, gender information, and other personally identifiable information. We present randomly selected data examples in Appendix A.

### v3

# JiraiBench: A Cross-lingual Benchmark for Evaluating Large Language Models’ Detection of Human Risky Health Behavior Content in Jirai Community

@@ Abstract
In this paper, we present the first cross-lingual dataset that captures a transnational cultural phenomenon, focusing on the Chinese and Japanese "Jirai" subculture and its association with risky health behaviors. Our dataset of more than 15,000 annotated social media posts forms the core of JiraiBench, a benchmark designed to evaluate LLMs on culturally specific content. This unique resource allowed us to uncover an unexpected cross-cultural transfer in which Japanese prompts better handle Chinese content, indicating that cultural context can be more influential than linguistic similarity. Further evidence suggests potential cross-lingual knowledge transfer in fine-tuned models. This work proves the indispensable role of developing culturally informed, cross-lingual datasets for creating effective content moderation tools that can protect vulnerable communities across linguistic borders.
JiraiBench: A Cross-lingual Benchmark for Evaluating Large Language Models’ Detection of Human Risky Health Behavior Content in Jirai Community
Yunze Xiao1†thanks: Equal Contribution §Project Leader Corresponding Authors , Tingyu He2†, Lionel Z. Wang3†§, Yiming Ma3, Xingyu Song4, Xiaohang Xu4, Mona Diab1‡ Irene Li4‡, Ka Chung Ng3‡ 1Carnegie Mellon University 2University of Washington 3The Hong Kong Polytechnic University 4The University of Tokyo {yunzex,mdiab}@cs.cmu.edu, zhe-leo.wang@connect.polyu.hk, ireneli@ds.itc.u-tokyo.ac.jp, kc-boris.ng@polyu.edu.hk
Disclaimer: This paper describes human content related to risky health behaviors and potentially harmful behaviors that may be disturbing to some readers.
@@ 1 Introduction
Risky health behaviors (RHB), such as eating disorders (ED), nonsuicidal self-injury (NSSI) and Drug Misuse (DM), constitute a profound public health challenge, often linked to complex mental health conditions (Baumeister and Scher, 1988; Van der Kolk et al., 1991; Firestone and Seiden, 1990). The digital era has complicated this landscape, as online communities can normalize or reinforce these behaviors, creating new challenges for intervention. What makes these behaviors particularly concerning is their tendency to operate beneath the surface of conventional detection systems, often interwoven with complex mental health conditions such as depression, anxiety, post-traumatic stress disorder, or personality disorders (Van der Kolk et al., 1991; Firestone and Seiden, 1990).
Previous research on detecting risky health behaviors online has established important methodological foundations but remains limited in real-world applicability. Studies have typically focused narrowly on isolated behaviors like eating disorders (ED) (Wang et al., 2017a; Moessner et al., 2018; Chancellor et al., 2016a), non-suicidal self-injury (NSSI) (Wang et al., 2017b; Un Nisa and Muhammad, 2021; Ragheb et al., 2021), or drug misuse (DM) (Fisher et al., 2023; Nasralah et al., 2020; Phan et al., 2017; Fan et al., 2017). These approaches address the complex comorbidities common in at-risk individuals. Furthermore, an overwhelming concentration on English content (Scherr et al., 2020; Tébar and Gopalan, 2021; Sixto-García and Duarte-Melo, 2020) creates significant blind spots in multilingual environments. While community-based detection has shown promise (Tébar and Gopalan, 2021; Wang et al., 2017a; Chancellor et al., 2016b), its limited adoption neglects the crucial, nuanced social dynamics through which these behaviors are normalized. These constraints highlight a collective need for more comprehensive, multilingual, and contextually-aware detection frameworks.
To address these gaps, our study examines the Jirai community (in English: landmine; in Chinese: 地雷; in Japanese: じらい) [^1] Jirai Kei, Aesthetics Wiki, a transnational phenomenon of social media that spans Chinese and Japanese online spaces. This community is an ideal case study as it simultaneously encompasses multiple RHBs (DM, ED, and NSSI), uses coded language to evade moderation, and operates across different regulatory environments. By analyzing this bilingual community, we demonstrate the necessity of incorporating language-specific nuances and community dynamics into detection systems, addressing the key limitations of existing models.
@@
  1. 1.
    We conducted comprehensive experiments in four state-of-the-art LLMs using two baseline configurations, systematically evaluating performance in both target languages with prompts in Chinese, Japanese, and English, revealing unexpected patterns of cross-cultural transfer.
  2. 2.
    We uncover a noteworthy emergent phenomenon in which Japanese instruction prompts consistently outperform Chinese prompts when processing Chinese content, suggesting important linguistic-cultural bridges in transnational content moderation tasks.
  3. 3.
    We investigate cross-lingual transfer capabilities by fine-tuning Qwen2.5 7B on Chinese data and demonstrate significant performance improvements on both source and target languages, particularly by improving detection of risky health behaviors in Japanese content without explicit Japanese training data.
@@ 2.1 Social Media and Mental Health
Natural language processing (NLP) has long been a pivotal tool for researchers who studied mental health problems through textual data. Social media are one of the valuable resources to understand the mental dynamics of users, with studies showing that online activities of individuals carry detectable signals and can reflect their psychological states (Paul and Dredze, 2011; Coppersmith et al., 2014).
@@
Among the considerable number of NLP-based mental health studies, depression is the most discussed because of its high prevalence and broad implications. It has been well established that the mental states of people can be conveyed through language use (Chancellor and De Choudhury, 2020) and that depressed people tend to use first-person pronouns and negative words with a higher frequency (De Choudhury et al., 2013). The boundary of NLP research in mental health was also extended to other types as mental illness has become a shared, concerning issue over the society. Post-Traumatic Stress Disorder (PTSD) (Murarka et al., 2021), Attention Deficit Hyperactivity Disorder (ADHD) (Guntuku et al., 2023), and other general problems have all been examined via language cues, such as references to trauma, heightened expressions of fear, expressions of worry, and rumination. These linguistic markers are emblematic of negative emotions, exposing critical insights into how mental states manifest in languages (Teodorescu et al., 2023).
@@
The evolution of NLP nurtured the growing research interest in this domain. As this field matured in mid-2010s, research focus shifted from domain-specific lexical analysis to machine learning-based predictive models. Behavioral features, such as linguistic style, posting frequency, and social network attributes were engineered and then fed into classifiers, such as logistic regression (LR) and support vector machine (SVM), to detect relevant user-generated content (De Choudhury et al., 2013; Tsugawa et al., 2015).
@@
The introduction of Transformer-based models revolutionized the research landscape in this area, bringing remarkable improvements in the detection task and outperformed earlier deep learning models (Jiang et al., 2020; Matero et al., 2019). In particular, specialized models like MentalBERT that were trained in a mental health-specific corpus have demonstrated an improved capability of identifying edged language patterns in people’s mental health disclosures (Ji et al., 2022).
@@
Although the prior methods have shown promising results, several key challenges persist. One recurring issue is data scarcity and quality. Clinical diagnostic data is normally private and scarce, while datasets populated with online posts were in the doubt of construct validity—the lack of a standardized assessment of each post not only introduces mislabeling risk but also complicates the process of results evaluation. Another pivotal issue lies in the peculiar, intricate linguistic patterns in posts involved with negative mental conditions. Usually, distress was expressed in a figurative way, where creative metaphors, sarcasm, and irony were communicated as an encoded language (Coppersmith et al., 2014; Mendes and Caseli, 2024). Furthermore, most of the research was developed on an English data set with far fewer in other languages (Cao et al., 2025) and the models therein were trained on a corpus constructed from a single platform or a homogeneous demographic, which consequently limits the applicability of the models to more diverse populations.
2.2 Risky Health Behaviors Detections
@@
Besides the shared attributes, targeting RHBs differs remarkably from the task on general mental illness. RHBs that might cause detrimental and even life-threatening outcomes necessitate more accurate detection to take preventive actions, while their unique linguistic patterns confront researchers with more challenges.
Among these challenges, the use of risky language is a prominent one. Individuals involved in RHBs tend to communicate their situations by using strategically encoded languages—expressions that are linguistically ambiguous, metaphorical, figurative, or culturally unique (Yadav et al., 2020; Mendes and Caseli, 2024). Therefore, effective RHBs scrutiny requires systems capable of reasoning over latent implications and culture-specific metaphors.
@@
Flagging users on social media who reveal a tendency towards NSSI and suicidal ideation is a vital yet sensitive application of NLP. Studies reveal that the combination of linguistic signals from content, such as hopelessness, self-hate, and past suicidal attempts, coupled with online behavioral patterns greatly improve the models performance—and it usually gets better when enhanced with domain knowledge (Zhang et al., 2024; Ng et al., 2023; Choudhury and Kıcıman, 2017). Some works focus specifically on NSSI that might include mentions of self-cutting or injury (Cliffe et al., 2021), where the incorporation of psychological lexicons helps models identify depressive expressions as warning sign.
@@
Likewise, ED detection is another intractable task because the involved individuals barely explicitly state their conditions. Instead, such tendencies manifest themselves through obsessive discussions of weight, food, calories, and body images, making ingestion of specific knowledge essential to ensure the performance of the models in the classification task (Chan et al., 2022).
@@
Social media-based DM study is also one concerning area (Wright et al., 2021). One unique challenge is the language ambiguity of drug references and the fast-evolving drug landscape, further complicating the construction of datasets—from identifying emerging drugs through obscure drug references to discerning the subtextual intent. Annotating such data also requires expert assistance in analyzing if the post is truly indicative of DM rather than casual innocuous mentioning.
The emergence of LLM has opened new research avenues, showing strong capabilities in detecting user-generated RHB content and identifying worsening psychological conditions, even in zero-shot or few-shot settings (Stade et al., 2024). These models excel at understanding linguistic nuances in intense RHB content, enabling them to detect subtextual signals of mental distress for immediate intervention. However, LLM-based RHB detection faces significant generalizability challenges across platforms, cultures, and languages. Since mainstream LLMs were trained primarily on English-centric corpora with limited representation of other languages, even state-of-the-art models struggle in multilingual environments where community-based, culture-dependent expressions are common. Additionally, while LLMs demonstrate strong contextual understanding, user-specific linguistic peculiarities can compromise performance, particularly when expressions are intentionally obfuscated to evade content moderation Xiao et al. (2024).
Jirai Kei "Jirai Kei" (in English: landmine style; in Chinese: 地雷系; in Japanese: じらいけい) refers to a transnational social media subculture that emerged in Japan during the early 2020s, gaining particular prominence during pandemic restrictions. The term literally translates to "landmine style", metaphorically representing psychological vulnerabilities that might "explode" when triggered. This community is characterized by individuals who publicly express various forms of psychological distress and self-destructive tendencies across social media platforms. Jirai communities exhibit distinctive communication patterns, including specialized terminology and coded language that conveys emotional vulnerability, particularly related to eating disorders, self-harm behaviors, and substance misuse. This phenomenon spans both Japanese and Chinese online spaces, creating interconnected networks where participants share experiences of psychological distress through culturally specific expressions.
@@ 3 Dataset Construction
In this section, we describe the construction of the JiraiBench dataset. We start with introducing the data collection and filtering process, then turning to the annotation.
@@ 3.1 Data Collection and Cleaning
For the Chinese dataset, we implemented a keyword-based search approach to gather relevant posts from Sina Weibo [^2] https://weibo.com/, China’s equivalent of X (formerly Twitter). Using specialized lexicons developed by community experts in both languages for DM, ED, and NSSI, we identified posts that discuss RHBs within the community. This process yielded a comprehensive Chinese dataset containing 10,419 potentially concerning posts. The lexicons were meticulously curated to include both explicit and implicit terminology commonly associated with each type of behavior.
Similarly, we applied the same methodological framework to construct the Japanese dataset, ensuring consistency between the collections. From X [^3] https://x.com/, we collected 5,000 tweets, with varying associations to the three behaviors of interest. Both datasets underwent an identical data cleansing pipeline to remove irrelevant content, including advertisements, duplicate entries, excessively short posts lacking meaningful information, and other noise. In accordance with the platform policies and to maintain user privacy, we eliminated all usernames, gender information, and other personally identifiable information. We present randomly selected data examples in Appendix B.